OpenStudy (anonymous):

what is the equation of the line in slope-intercept form that passes through the point (1,3) and is perpendicular to the line y=-x+2?

OpenStudy (anonymous):

perpendicular lines have opposite of reciprocal slopes in other words if u multiply the slopes u should get negative 1 say m is the slope of the line u want m(-1)=-1 so m=1 y=x+b now solve with the point u have 3=1+b 2=b y=x+2